# Synthetic-to-Real Camouflaged Object Detection

### v1

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755461isbn: 979-8-4007-2035-2/2025/10conference: Make sure to enter the correct conference title from your rights confirmation email; October 27–31, 2025; Dublin, Irelandccs: Computing methodologies Scene understandingccs: Computing methodologies Image segmentation
@@ 2.2. Unsupervised Domain Adaptation
The core idea of Unsupervised Domain Adaptation (UDA) is to improve performance on the target domain by learning from the source domain and narrowing the distributional gap between the source and target domain without requiring labeled data from the target domain. Among the various techniques developed for domain adaptation, three commonly explored approaches include divergence-based methods, adversarial learning, and image-to-image translation with style transfer. a) Divergence: MMD (Gretton et al., 2006, 2012) is a statistical test used to determine if two distributions are equivalent based on their samples. CAN (Kang et al., 2019) minimizes cross-entropy loss on labeled target data while alternating between estimating labels for target samples. b) Adversarial Learning: DANN (Ajakan et al., 2014; Ganin and Lempitsky, 2015; Ganin et al., 2016) train the feature extractor by negating the gradient from the domain classifier using a gradient reversal layer during backpropagation. In ADDA (Tzeng et al., 2017), the feature extractor is trained by inverting the labels for the domain classifier, making it learn domain-invariant representations. Shen et al. (Shen et al., 2018) proposed WDGRL by modifying DANN, replacing the domain classifier with a network that directly approximates the Wasserstein distance. Recently, teacher-student paradigms (Tarvainen and Valpola, 2017) such as PMT (Belal et al., 2024) have been explored to enhance target representation learning through soft supervision and class-level alignment. C) Image-to-Image Translation and Style Transfer: CycleGAN (Zhu et al., 2017), based on pix2pix, is a commonly used unsupervised image-to-image translation method, along with similar approaches like DualGAN (Yi et al., 2017) and DiscoGAN (Kim et al., 2017). FDA (Yang and Soatto, 2020) uses Fast Fourier Transform to propose a simple method for domain alignment that does not require any learning. In addition to these classical techniques, recent efforts have started leveraging large-scale pre-trained vision-language models to improve cross-domain generalization (Lai et al., 2024a). Unlike the aforementioned methods, our approach explicitly frames UDA as a robust and efficient noisy label learning task, simplifying the challenge and delivering enhanced performance.
@@ 3.2. Stage A: Inter-Domain Transfer
$\min\limits_{f}$ $\mathbb{E}_{(x_{s},y_{s})\sim\mathcal{D}_{s},x_{t}\sim\mathcal{D}_{t}}\bigl{[}\mathcal{L}_{CE}\bigl{(}f\bigl{(}\mathcal{A}\bigl{(}x_{s}\bigr{)}\bigr{)},y_{s}\bigr{)}\bigr{]}$
@@ 4.1. Settings
[…]
@@ 4.2. Results and Analysis
Comparison with Baselines. In Tab. 1 and Tab. 2, we compare the Source-Only baseline, Mean Teacher (Tarvainen and Valpola, 2017), and our method on $\hat{S}2C$ and $\hat{C}2C$ tasks. From the experimental results, we observe that Source-Only exhibits limited performance across all models, indicating a significant domain gap between synthetic and real images. Mean Teacher improves model performance to some extent by introducing target domain data. However, since this method does not filter the pseudo labels generated by the teacher model, a performance drop is observed on certain metrics for some models. For example, in $\hat{S}2C$ task, both SINet (Fan et al., 2020) and SINet-v2 (Fan et al., 2021a) under the Mean Teacher method show a certain decrease in MAE and $E_{\phi}^{ad}$ compared to the Source-Only baseline, with the $E_{\phi}^{ad}$ metric averaging a 2% drop. This suggests that unfiltered pseudo labels may introduce noise, which negatively impacts the model performance. Our proposed method outperforms both Source-Only and Mean Teacher on all metrics in $\hat{S}2C$ and $\hat{C}2C$ tasks. For example, on SINet, $F_{\beta}^{w}$ increased by 11.1% and 12%, and $F_{\beta}^{mx}$ increased by 12.5% and 13.5%. Compared to Mean Teacher, $F_{\beta}^{w}$ and $F_{\beta}^{mx}$ also increased by 6.4%, 5%, and 3.6%, 4.4% respectively. By comparing the overall experimental data in Tab. 1 and Tab. 2, it can be observed that the performance on $\hat{S}2C$ task is significantly lower than that on $\hat{C}2C$ tasks across different models. This result indicates that, despite using synthetic COD datasets as the source domain, there is still a significant domain gap between SOD and COD. However, it is worth noting that despite this domain gap, synthesizing a large amount of COD data from easily accessible SOD images remains a valuable research direction. Synthetic data provides an initial training basis for the model, especially in scenarios where the cost of labeling real COD data is high (Chen et al., 2024; Li et al., 2022; Sun et al., 2022; Zhuge et al., 2022; Zhu et al., 2022), and can effectively alleviate the problem of data scarcity. As documented in Tab. 3, we conduct comprehensive evaluations on three mainstream COD datasets using SINet as the base model, validating the effectiveness of our proposed method. Compared with Source-Only and Mean Teacher, the proposed method presents a substantial advantage in terms of key performance indicators.
@@
[…]

### v3

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755461isbn: 979-8-4007-2035-2/2025/10conference: Make sure to enter the correct conference title from your rights confirmation email; October 27–31, 2025; Dublin, Irelandccs: Computing methodologies Scene understandingccs: Computing methodologies Image segmentation
@@ 2.2. Unsupervised Domain Adaptation
The core idea of Unsupervised Domain Adaptation (UDA) is to improve performance on the target domain by learning from the source domain and narrowing the distributional gap between the source and target domain without requiring labeled data from the target domain. Among the various techniques developed for domain adaptation, three commonly explored approaches include divergence-based methods, adversarial learning, pseudo labeling and image-to-image translation with style transfer. a) Divergence: MMD (Gretton et al., 2006, 2012) is a statistical test used to determine if two distributions are equivalent based on their samples. CAN (Kang et al., 2019) minimizes cross-entropy loss on labeled target data while alternating between estimating labels for target samples. b) Adversarial Learning: DANN (Ajakan et al., 2014; Ganin and Lempitsky, 2015; Ganin et al., 2016) train the feature extractor by negating the gradient from the domain classifier using a gradient reversal layer during backpropagation. In ADDA (Tzeng et al., 2017), the feature extractor is trained by inverting the labels for the domain classifier, making it learn domain-invariant representations. Shen et al. (Shen et al., 2018) proposed WDGRL by modifying DANN, replacing the domain classifier with a network that directly approximates the Wasserstein distance. c) Pseudo Labeling: Recently, teacher-student paradigms (Tarvainen and Valpola, 2017) have been introduced to enable self-training for UDA, especially the source-free UDA (SFDA) task. For example, SSNLL (Chen et al., 2022) formulate SFDA task as a noisy label learning problem, and employ the teacher-student to tackle this problem. And some works (Liu et al., [n. d.]; Lin et al., 2023) focus on the cooperation between the teacher and student models. d) Image-to-Image Translation and Style Transfer: CycleGAN (Zhu et al., 2017), based on pix2pix, is a commonly used unsupervised image-to-image translation method, along with similar approaches like DualGAN (Yi et al., 2017) and DiscoGAN (Kim et al., 2017). FDA (Yang and Soatto, 2020) uses Fast Fourier Transform to propose a simple method for domain alignment that does not require any learning. In addition to these classical techniques, recent efforts have started leveraging large-scale pre-trained vision-language models to improve cross-domain adaptation (Lai et al., 2024a; Chen et al., 2025). Unlike the aforementioned methods, our approach explicitly frames UDA as a robust and efficient noisy label learning task, simplifying the challenge and delivering enhanced performance.
@@ 3.2. Stage A: Inter-Domain Transfer
$\min\limits_{f}\$$\mathbb{E}_{(x_{s},y_{s})\sim\mathcal{D}_{s},x_{t}\sim\mathcal{D}_{t}}\bigl[\mathcal{L}_{CE}\bigl(f\bigl(\mathcal{A}\bigl(x_{s}\bigr)\bigr),y_{s}\bigr)\bigr]$
@@ 4.1. Settings
Evaluation Metrics. Referring to previous work (Zhang and Wu, 2023; Lai et al., 2024b; Hu et al., 2024; Yang et al., 2021; He et al., 2023a), we use five metrics to comprehensively evaluate the model performance in COD task, including structure measure $(S_{\alpha})$ (Fan et al., 2017), weighted F-measure $(F_{\beta}^{w})$ (Margolin et al., 2014), enhanced-alignment measure $(E_{\varphi})$ (Fan et al., 2018, 2021b), F-measure $(F_{\beta})$ (Borji et al., 2015), and mean absolute error $(M)$.
@@ 4.2. Results and Analysis
Comparison with Baselines. In Tab. 1 and Tab. 2, we compare the Source-Only baseline, Mean Teacher (Tarvainen and Valpola, 2017), and our method on $\hat{S}2C$ and $\hat{C}2C$ tasks. From the experimental results, we observe that Source-Only exhibits limited performance across all models, indicating a significant domain gap between synthetic and real images. Mean Teacher improves model performance to some extent by introducing target domain data. However, since this method does not filter the pseudo labels generated by the teacher model, a performance drop is observed on certain metrics for some models. For example, in $\hat{S}2C$ task, both SINet (Fan et al., 2020) and SINet-v2 (Fan et al., 2021a) under the Mean Teacher method show a certain decrease in MAE and $E_{\phi}^{ad}$ compared to the Source-Only baseline, with the $E_{\phi}^{ad}$ metric averaging a 2% drop. This suggests that unfiltered pseudo labels may introduce noise, which negatively impacts the model performance. Our proposed method outperforms both Source-Only and Mean Teacher on all metrics in $\hat{S}2C$ and $\hat{C}2C$ tasks. For example, on SINet, $F_{\beta}^{w}$ increased by 11.1% and 12%, and $F_{\beta}^{mx}$ increased by 12.5% and 13.5%. Compared to Mean Teacher, $F_{\beta}^{w}$ and $F_{\beta}^{mx}$ also increased by 6.4%, 5%, and 3.6%, 4.4% respectively. By comparing the overall experimental data in Tab. 1 and Tab. 2, it can be observed that the performance on $\hat{S}2C$ task is significantly lower than that on $\hat{C}2C$ tasks, highlighting a substantial domain gap between SOD and COD. However, it is worth noting that despite this domain gap, synthesizing a large amount of COD data from easily accessible SOD images remains a valuable research direction. Synthetic data provides an initial training basis for the model, especially in scenarios where the cost of labeling real COD data is high (Chen et al., 2024; Li et al., 2022; Sun et al., 2022; Zhuge et al., 2022; Zhu et al., 2022), and can effectively alleviate the problem of data scarcity. As documented in Tab. 3, we conduct comprehensive evaluations on three mainstream COD datasets using SINet as the base model, validating the effectiveness of our proposed method. Compared with Source-Only and Mean Teacher, the proposed method presents a substantial advantage in terms of key performance indicators.
@@
Figure 4. The impact of different hyperparameter combinations $\alpha$, $\beta$, $\delta$ on ES Loss under the HKU-IS $\rightarrow$ COD10K($\hat{S}2C$) and CAMO + NC4K + CHAM. $\rightarrow$ COD10K($\hat{C}2C$) benchmarks. The left figure presents the model performance in terms of $S_{\alpha}$, while the right figure shows the results for $E_{\phi}^{mn}$.